Grade 11Algebra

How to solve this inequality : x2+1>0?
Have a problem solving this one. Please help me

Nakul Jadeja
9 Years agoGrade 11

1 Answer

Dinesh Nayak
9 Years ago
for the given inequality x2+1>0
 
on substituting the values for x such as :- { -\infty ,,,,,,,-30,,,,, -25,,,, -15,,,, -9,.... -2, -1, 0, 1,2,..... 7, …..12,....25.....60......\infty}
 
so  the solution set of the inequality = ( – \infty , \infty )